# RxRx3-core: Benchmarking drug-target interactions in High-Content Microscopy

### v1

@@ Abstract
High Content Screening (HCS) microscopy datasets have transformed the ability to profile cellular responses to genetic and chemical perturbations, enabling cell-based inference of drug-target interactions (DTI). However, the adoption of representation learning methods for HCS data has been hindered by the lack of accessible datasets and robust benchmarks. To address this gap, we present RxRx3-core, a curated and compressed subset of the RxRx3 dataset, and an associated DTI benchmarking task. At just 18GB, RxRx3-core significantly reduces the size barrier associated with large-scale HCS datasets while preserving critical data necessary for benchmarking representation learning models against a zero-shot DTI prediction task. RxRx3-core includes 222,601 microscopy images spanning 736 CRISPR knockouts and 1,674 compounds at 8 concentrations. RxRx3-core is available on HuggingFace and Polaris, along with pre-trained embeddings and benchmarking code, ensuring accessibility for the research community. By providing a compact dataset and robust benchmarks, we aim to accelerate innovation in representation learning methods for HCS data and support the discovery of novel biological insights.
@@ 1 Introduction
To address these limitations, we introduce RxRx3-core, a dataset specifically designed to facilitate research in representation learning for HCS data. At only 18 GB (17.5GB of images and 0.5Gb of pre-computed embeddings), RxRx3-core is significantly more accessible than its predecessors and is accompanied by benchmarks evaluating perturbation signal magnitude and zero-shot DTI prediction. By providing both a manageable dataset and well-defined benchmarks, we aim to accelerate innovation in representation learning methods for high-content phenotypic screening.
@@ 4 Results
RxRx3-core enables the computation of benchmarks that evaluate the quality of perturbation embeddings in capturing biological signals across key tasks: distinguishing perturbations from control populations and zero-shot retrieval of drug-target interactions. Before calculating these benchmarks, crop aggregation and batch alignment techniques are applied to phenomics embeddings from representation learning models. First, the four tiled embeddings from every 256x256x6 crop of each 512x512x6 microscopy image are mean-aggregated. Next, a crucial batch alignment step is used to center the latent space on control samples and align experimental batches, ensuring comparability across conditions (Celik et al., 2024).
@@ 4.2 Drug-target interaction benchmark
5 Conclusion
We introduce RxRx3-core, a compact and accessible dataset with an associated set of benchmarks designed to evaluate representation learning models for high-content screening microscopy data against zero-shot drug-target interaction prediction. At just 18GB, RxRx3-core is available on HuggingFace and Polaris, making it practical for widespread use. Our analysis includes baseline and foundation model performance across perturbation signal magnitude and zero-shot drug-target interaction prediction benchmarks. The results highlight the advantages of self-supervised vision transformer models, particularly as they scale to multi-billion parameter regimes, in capturing biologically meaningful representations. We hope that RxRx3-core and its benchmarks serve as a valuable resource for the machine learning community, fostering innovation in representation learning approaches and advancing the analysis of large-scale microscopy datasets.
Meaningfulness Statement
Acknowledgments
Appendix A Appendix
A.1 Gene-gene interaction benchmark
RxRx3-core can potentially be used to benchmark representation learning models against zero-shot gene-gene interaction prediction. However, we caution users from making conclusions based on evaluating this task on the RxRx3-core dataset given the relatively small number of CRISPR knockouts in the dataset (736). The task, previously described in Celik et al. (2024) and evaluated in Kraus et al. (2024) and Kenyon-Dean et al. (2024), leverages embeddings from representation learning models to build biological maps that can be used to uncover novel and meaningful interactions between perturbations. To assess the ability of a map to represent known biological relationships, we employ four annotation sources containing gene-gene interactions as benchmarks (Table 5).
@@ A.1 Gene-gene interaction benchmark
CORUM Giurgiu et al. (2019) Gene clusters corresponding to protein complexes
HuMAP (Drew et al., 2017) Gene clusters representing protein complexes
@@
The core hypothesis is that a map capable of accurately capturing known relationships provides a strong representation of biological systems and is likely to reveal novel interactions. To evaluate this, closely following Celik et al. (2024) Multivariate known biological relationship recall benchmark, we compute recall for the most extreme 10% of pairwise relationships based on cosine similarity between embeddings. This includes both the top 5% (high similarity) and bottom 5% (low similarity) of the pairwise similarity distribution, as the latter may indicate opposing functions between genes. A random or uninformative map serves as a baseline, expected to achieve a recall of 10%.
To calculate recall, we first compute pairwise cosine similarities between the aggregated embeddings of all perturbed genes in RxRx3-core. Self-links, where a gene is compared to itself, are excluded because their cosine similarity is always one, which would bias the results. “Predicted links” are defined as relationships falling within the top or bottom 5% of the similarity distribution. Recall is then determined for each annotation source as the proportion of true predicted links relative to all possible links. Since datasets differ in the number of genes they include, we adjust recall calculations to consider only gene-gene interactions where both genes are present in the dataset. Therefore, recall results (Table 6) from the subset of 736 gene knockouts (KOs) in RxRx3-core are not directly comparable with other work (Kraus et al., 2024; Kenyon-Dean et al., 2024) which measures recall across a broader set of 17,000 genes.
To demonstrate how models rank on this metric when a larger set of gene-KOs are leveraged, we report the same gene-gene relationship recall benchmark (Celik et al., 2024) on the full RxRx3 dataset, including 17,063 gene-KOs, and on the JUMP dataset (cpg0016) (Chandrasekaran et al., 2023), including a subset of 7,976 gene-KOs, in Table 6. As shown in the table, the number of interactions evaluated grows significantly as more CRISPR knockouts are included in the datasets. For post-processing RxRx3-core and JUMP-CP embeddings, we use PCA with center-scaling for standard batch correction alignment. We observe that the MAEs perform better than the CellProfiler feature extraction baseline (Carpenter et al., 2006; Kamentsky et al., 2011) when a larger set of genes and gene-gene interactions are evaluated. For post-processing the full RxRx3 dataset, we correct for batch effects using Typical Variation Normalization (TVN) (Ando et al., 2017), and also correct for possible chromosome arm biases known to exist in CRISPR-Cas9 HCS data (Lazar et al., 2024). Notably, recall on JUMP-CP is considerably lower than on RxRx3 likely due to different assay protocols and more variance in the data.
Model CORUM HuMAP Reactome StringDB
@@
CellProfiler .748 .813 .473 .652
OpenPhenom-S/16 .654 .709 .436 .579
Phenom-1 .721 .784 .402 .623
Phenom-2 .731 .810 .407 .647
@@
Table 6: Multivariate known biological relationship recall estimated @ 0.05-0.95 percentiles for benchmark databases (CORUM, hu.MAP, Reactome-PPI, and StringDB) over 3 random seeds sampling the null distribution (SD ±$\leq\pm 10^{-3}$). For RxRx3-core dataset the embeddings are batch-corrected using a PCA-CS aligner. For RxRx3 dataset, the embeddings are transformed with a TVN aligner Ando et al. (2017) followed by a chromosome arm bias correction Lazar et al. (2024). For JUMP-CP (cpg0016) dataset we used a PCA-CS aligner.

### v2

@@ Abstract
High Content Screening (HCS) microscopy datasets have transformed the ability to profile cellular responses to genetic and chemical perturbations, enabling cell-based inference of drug-target and gene-gene interactions. However, the adoption of representation learning methods for HCS data has been hindered by the lack of accessible datasets and robust benchmarks. To address this gap, we present RxRx3-core, a curated and compressed subset of the RxRx3 dataset, and associated drug-target and gene-gene interaction benchmarks. At just 18GB, RxRx3-core significantly reduces the size barrier associated with large-scale HCS datasets while preserving critical data necessary for benchmarking representation learning models against zero-shot perturbation interaction prediction tasks. RxRx3-core includes 222,601 microscopy images spanning 736 CRISPR knockouts and 1,674 compounds at 8 concentrations. RxRx3-core is available on HuggingFace and Polaris, along with pre-trained embeddings and benchmarking code, ensuring accessibility for the research community. By providing a compact dataset and robust benchmarks, we aim to accelerate innovation in representation learning methods for HCS data and support the discovery of novel biological insights.
@@ 1 Introduction
To address these limitations, we introduce RxRx3-core, a dataset specifically designed to facilitate research in representation learning for HCS data. At only 18 GB (17.5GB of images and 0.5Gb of pre-computed embeddings), RxRx3-core is significantly more accessible than its predecessors and is accompanied by benchmarks evaluating perturbation signal magnitude and zero-shot DTI and gene-gene interaction prediction. By providing both a manageable dataset and well-defined benchmarks, we aim to accelerate innovation in representation learning methods for high-content phenotypic screening.
@@ 4 Results
RxRx3-core enables the computation of benchmarks that evaluate the quality of perturbation embeddings in capturing biological signals across key tasks: distinguishing perturbations from control populations and zero-shot retrieval of drug-target and gene-gene interactions. Before calculating these benchmarks, crop aggregation and batch alignment techniques are applied to phenomics embeddings from representation learning models. First, the four tiled embeddings from every 256x256x6 crop of each 512x512x6 microscopy image are mean-aggregated. Next, a crucial batch alignment step is used to center the latent space on control samples and align experimental batches, ensuring comparability across conditions (Celik et al., 2024).
@@ 4.2 Drug-target interaction benchmark
4.3 Gene-gene interaction benchmark
In addition to benchmarking drug-target interactions, RxRx3-core can be used to benchmark representation learning models against zero-shot gene-gene interaction prediction. The task, previously described in Celik et al. (2024) and evaluated in Kraus et al. (2024) and Kenyon-Dean et al. (2024), leverages embeddings from representation learning models to build biological maps that can be used to uncover novel and meaningful interactions between perturbations. To assess the ability of a map to represent known biological relationships, we employ four annotation sources containing gene-gene interactions as benchmarks (Table 5).
@@ 4.3 Gene-gene interaction benchmark
CORUM Giurgiu et al. (2019) Highly-curated gene clusters of protein complexes
hu.MAP (Drew et al., 2017) Comprehensive atlas of interactions and protein complexes
@@
The core hypothesis is that a map capable of accurately capturing known relationships provides a strong representation of biological systems and is likely to reveal novel interactions. To evaluate this, closely following Celik et al. (2024) Multivariate known biological relationship recall benchmark, we compute recall for the most extreme 10% of pairwise relationships based on cosine similarity between embeddings. This includes both the top 5% (high similarity) and bottom 5% (low similarity) of the pairwise similarity distribution, as the latter may indicate opposing functions between genes; for example, we often find that the gene knockouts for MTOR and TSC2 point in opposite directions with significant negative cosine similarities to each other. A random or uninformative map serves as a baseline, expected to achieve a recall of 10%.
Table 6: Multivariate known biological relationship recall obtained below the 5th and above the 95th percentiles (0.05-0.95 threshold) of all gene-gene cosine similarities against the null distribution, as per Celik et al. (2024); Kraus et al. (2024); Kenyon-Dean et al. (2024). We report the mean recall over 3 different random seeds of sampling the null distribution in each benchmark run ($\sigma\leq\pm 0.001$), computed over four different gene-gene interaction benchmark databases: CORUM, hu.MAP, Reactome-PPI, and StringDB. For RxRx3-core and the JUMP-CP (cpg0016) the embeddings are batch-corrected using a PCA-CS aligner. For RxRx3 dataset, the embeddings are transformed with a TVN aligner (Ando et al., 2017) followed by a chromosome arm bias correction Lazar et al. (2024).
Model CORUM hu.MAP Reactome StringDB
@@
CellProfiler .361 .444 .160 .330
OpenPhenom-S/16 .300 .352 .158 .281
Phenom-1 .395 .482 .188 .349
Phenom-2 .486 .553 .197 .415
@@
To calculate recall, we first compute pairwise cosine similarities between the aggregated embeddings of all perturbed genes in RxRx3-core. Self-links, where a gene is compared to itself, are excluded because their cosine similarity is always one, which would bias the results. “Predicted links” are defined as relationships falling within the top or bottom 5% of the similarity distribution. Recall is then determined for each annotation source as the proportion of true predicted links relative to all possible links. Since datasets differ in the number of genes they include, we adjust recall calculations to consider only gene-gene interactions where both genes are present in the dataset. Therefore, recall results from the subset of 736 gene knockouts (KOs) in RxRx3-core are not directly comparable with other work (Kraus et al., 2024; Kenyon-Dean et al., 2024) which measures recall across a broader set of 17,000 genes. Despite the subset of genes available in RxRx3-core, Table 6 shows that self-supervised MAEs outperform CellProfiler features as the models are scaled to larger datasets and architectures.
To demonstrate how models rank on this metric when a larger set of gene-KOs are leveraged, we report the same gene-gene relationship recall benchmark (Celik et al., 2024) on the full RxRx3 dataset, including 17,063 gene-KOs, and on the JUMP dataset (cpg0016) (Chandrasekaran et al., 2023), including a subset of 7,976 gene-KOs, in Table 6. As shown in the table, the number of interactions evaluated grows significantly as more CRISPR knockouts are included in the datasets. For post-processing RxRx3-core and JUMP-CP embeddings, we use PCA with center-scaling for standard batch correction alignment. We observe that all MAEs, including OpenPhenom-S/16, perform better than the CellProfiler feature extraction baseline (Carpenter et al., 2006; Kamentsky et al., 2011) when a larger set of genes and gene-gene interactions are evaluated. For post-processing the full RxRx3 dataset, we correct for batch effects using Typical Variation Normalization (TVN) (Ando et al., 2017), and also correct for possible chromosome arm biases known to exist in CRISPR-Cas9 HCS data (Lazar et al., 2024). Notably, recall on JUMP-CP is considerably lower than on RxRx3 likely due to different assay protocols and more variance in the data.
5 Conclusion
We introduce RxRx3-core, a compact and accessible dataset with an associated set of benchmarks designed to evaluate representation learning models for high-content screening microscopy data against zero-shot drug-target and gene-gene interaction prediction. At just 18GB, RxRx3-core is available on HuggingFace and Polaris, making it practical for widespread use. Our analysis includes baseline and foundation model performance across perturbation signal magnitude and zero-shot drug-target and gene-gene interaction prediction benchmarks. The results highlight the advantages of self-supervised vision transformer models, particularly as they scale to multi-billion parameter regimes, in capturing biologically meaningful representations. We hope that RxRx3-core and its benchmarks serve as a valuable resource for the machine learning community, fostering innovation in representation learning approaches and advancing the analysis of large-scale microscopy datasets.
Meaningfulness Statement
Acknowledgments
Appendix A Appendix